OpenStudy (anonymous):

A copper alloy that is 40% copper is to be combined with a copper alloy that is 80% copper to produce 120 kilograms of an alloy that is 70% copper. How many kilograms of each alloy must be used?

OpenStudy (anonymous):

we need 70% of 120 kgs = 84 kgs copper 0.4x + 0.8y = 84 where x and y are weights of two alloys x + y = 120 solve to get the answer

OpenStudy (anonymous):

haha cant you just give me the answer :P

OpenStudy (anonymous):

jimmy i got x equals 30 and y eqyuals 90 is that right jimmy boy? :p

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

30 * 0.4 = 12 kgs 90*0.8 = 72 kgs totals 12 + 72 =84 30 + 90 = 120

OpenStudy (anonymous):

dead right